OpenStudy (anonymous):

5^2x+1=3^4x-1 Find x.

OpenStudy (anonymous):

\[5^{2x+1}=3^{4x-1}\]

OpenStudy (janu16):

Is x-1 included as a square root? I don't think so but just asking

OpenStudy (anonymous):

It is, sadly - or else it would be much easier. I wasn't sure how to take logs of each side.

OpenStudy (anonymous):

It's part of the power.

OpenStudy (janu16):

Let me try

OpenStudy (janu16):

I think its 2.3

OpenStudy (anonymous):

Take log of each side is good start. HINT: \(\log(a^n) = n\log(a)\) So you should end up with \((2x+1)\ln(5) = (4x+1)\ln(3)\)

OpenStudy (anonymous):

ln(5) and ln(3) are just constants, so now it's simple algebra. make sense so far?

OpenStudy (anonymous):

oops 4x-1* but idea is same

OpenStudy (anonymous):

final answer is going to be ugly lol...

OpenStudy (janu16):

@Vaudriel i have the steps if you want

OpenStudy (janu16):

I can show you steps so you will get it

OpenStudy (anonymous):

Would you please? Thanks

OpenStudy (anonymous):

What would you do after moving the exponents to the front of the terms?

OpenStudy (anonymous):

Would you take ln of 5 and multiply it with (2x+1)?

Nnesha (nnesha):

yes take ln of 5

Nnesha (nnesha):

and then distribute parentheses

OpenStudy (anonymous):

Okay, thanks!

Nnesha (nnesha):

got it ?? :O

OpenStudy (anonymous):

Yeah, so I combine like terms and divide?